OpenStudy (anonymous):

3. How many solutions are there to the following system of equations? 3x + 7y = –10 15x + 35y = –45 (Points : 4) 0 1 2 infinitely many

OpenStudy (agent0smith):

3x + 7y = –10 15x + 35y = –45 Can you see a number that you could multiply 3x by, to make it the same as 15x?

OpenStudy (anonymous):

is the answer 2

OpenStudy (agent0smith):

Show your work.

OpenStudy (agent0smith):

3x + 7y = –10 15x + 35y = –45 If you wanted to eliminate x, you could multiply the first equation by what number...?

OpenStudy (anonymous):

The equations are both of straight lines. When we are solving a system, we are trying to find out where they cross each other. There are 3 possibilities. Two lines can be parallel and they would intersect 0 times. (they'd never cross each other.) Two lines that are different could intersect each other at one point. The last possibility is that the two lines are the same and are basically on top of each other. We say they intersect at infinitely many points. If you divide the second equation by 5, you almost have the first equation. 3x + 7y = –10 3x + 7y = -9. Let's re-write them in a slope & y-intercept form. Just rearrange them and solve for y. y = (-3/7)x - 10/7 y = (-3/7)x - 9/7 These two lines have the same slope so they're parallel. They'll never meet. So the answer is 0. If you have any questions, please ask. Learning the steps will make your school work easier in the future.

OpenStudy (anonymous):

i just need the answer ppl

OpenStudy (agent0smith):

It was posted above if you looked carefully.